# Parity distributions among gaps of free numerical semigroups

### v1

@@ Abstract
In this paper, we extend recent results about the distribution of even and odd gaps of a numerical semigroup. We find that, for any numerical semigroup, the distribution can be computed in terms of the numbers of or the sums of odd and even elements in a corresponding Apéry set. With free numerical semigroups specifically, we show that there are always at least as many odd gaps as even gaps, with equality precisely when the generating elements are all odd. We then specialize these results to the cases of numerical semigroups generated by compound sequences.
@@ 1 Introduction
In [1], Cho, Lee, & Nam investigated the distribution of odd and even gaps of numerical semigroups. They gave general results in the case where one knows the Apéry set of a numerical semigroup relative to an odd generating element. For this, one needs only to know the difference between the numbers of even and odd elements of the Apéry set. They also gave explicit results for a few families of numerical semigroups which have Apéry sets that are known in the literature. Notably, in the case of a numerical semigroup $S$ generated by a compound sequence for which all minimal generating elements are odd, they showed that $S$ has as many odd gaps as it has even gaps. In their concluding remarks, they mentioned an open problem about computing the distribution of odd and even gaps in the case of a numerical semigroup generated by a compound sequence where at least one of the minimal generating elements is even.
@@ 1.1 Organization
In Section 3, we look specifically at the case of free numerical semigroups along with the sequences, known as telescopic sequences, that one uses to generate them. With knowledge of the Apéry sets of free numerical semigroups, we use an identity from Gassert & Shor [3] to compute the difference between the numbers of odd gaps and even gaps of any free numerical semigroup. Our main result is Theorem 3.5. Immediately afterward, we have two corollaries: with one we give a simple formula for the difference in the case where all but one of the generating elements of a free numerical semigroup are even; and with the other we show that a free numerical semigroup $S$ has the same number of odd gaps as even gaps precisely when the minimal generating elements of $S$ are all odd. We then apply these results to numerical semigroups generated by compound sequences and geometric sequences, and we round off this section with some explicit examples.
@@ 2.1 Preliminaries
For a comprehensive introduction to numerical semigroups, we refer the reader to [5].
@@
It is known that $\langle n_{1},\dots,n_{k}\rangle$ is a numerical semigroup if and only if $\gcd(n_{1},\dots,n_{k})=1$. Furthermore, every numerical semigroup $S$ can be written in the form $S=\langle n_{1},\dots,n_{k}\rangle$ for some sequence of positive integers $(n_{1},\dots,n_{k})$ with $\gcd(n_{1},\dots,n_{k})=1$. We say that $S$ is minimally generated by $(n_{1},\dots,n_{k})$ if no proper subsequence generates $S$. It happens every numerical semigroup $S$ has a unique minimal generating sequence. Such a minimal generating sequence is necessarily finite.
@@
Put another way, $\operatorname{Ap}(S;t)$ consists of exactly $t$ elements, each the smallest element of $S$ in its congruence class modulo $t$. Many properties of a numerical semigroup can be derived directly from one of its Apéry sets. We will use the following identity from Gassert & Shor [3].
Proposition 2.2 ([3, Theorem 2.3]).
@@
where $\sigma(T)$ denotes the sum of the elements of a finite set $T$. (This result is originally due to Selmer. See [6, Section 2].)
@@ 2.2 Distribution of odd and even gaps given any Apéry set
Cho, Lee, & Nam give a formula ([1, Theorem 3.1]) for O(G(S))E(G(S))O(G(S))-E(G(S)) in terms of $\operatorname{Ap}(S;t)$ in the case where $t$ is an odd minimal generating element of $S$. With Proposition 2.2, we are able to extend the result for any nonzero $t\in S$, odd or even. We also note that $t$ need not be a generating element of $S$ here.
@@
If $t$ is even, we use $f(n)=n(-1)^{n}$ with Proposition 2.2.
@@
Since $t$ is even, $(-1)^{t+g}=(-1)^{g}$ and $\sum_{k=0}^{t-1}=k(-1)^{k}=-t/2$. Thus,
@@
consisting of 16 odd gaps and 9 even gaps. We can verify directly that O(G(S))=E(G(S))=169=7O(G(S))=E(G(S))=16-9=7.
@@ 3.1 Telescopic sequences and free numerical semigroups
We are now ready to specialize our results to free numerical semigroups. For a more detailed treatment, see [7] and the references therein.
@@
There are many families of telescopic sequences. For instance, any two-term $(a,b)$ sequence is telescopic, as is any geometric sequence $(a^{k},a^{k-1}b,\dots,b^{k})$. As we will see Section 3.3, compound sequences are telescopic as well.
@@ 3.2 Distribution of odd and even gaps of a free numerical semigroup
$F_{j}=\prod\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{k}\sum\limits_{n_{i}=0}^{c_{i}-1}(-1)^{n_{i}t_{i}}\sum\limits_{n_{j}=0}^{c_{j}-1}n_{j}t_{j}(-1)^{n_{j}t_{j}}.$
@@
Now, suppose $m>0$, which implies $\{0,\dots,m-1\}\subseteq I$. We need to show that
@@ 3.3 Specialization to compound sequences
We now focus on numerical semigroups generated by compound sequences, which were first described in [4].
@@
Note that the definition of a compound sequence in [4] specifies that $2\leq a_{i}<b_{i}$ for all $i$. With this condition in place, one has that $(n_{0},\dots,n_{p})$ is an increasing sequence. This is unnecessary for our work so we make no such restriction. The condition that $\gcd(a_{i},b_{j})=1$ for all $i\geq j$ is sufficient for our purposes.
The following proposition, from [2], says that the condition where $a_{i},b_{i}\geq 2$ for all $i$ gives rise to a minimal generating sequence.
@@
[2, Propositions 2.6 and 2.7] Suppose $(A,B)$ is a suitable pair with $C(A,B)=(n_{0},\dots,n_{p})$. Then $\gcd(n_{0},\dots,n_{p})=1$, implying that $\langle n_{0},\dots,n_{p}\rangle$ is a numerical semigroup. If we further have that $a_{i},b_{i}\geq 2$ for all $i$, then $(n_{0},\dots,n_{p})$ is a minimal generating sequence for this numerical semigroup.
@@
Proposition 3.10 ([3, Corollary 3.11]).
If $CS=(n_{0},\dots,n_{p})$ is a compound sequence, then $CS$ is a telescopic sequence. In particular, if $CS=C(A,B)$ for a suitable pair $(A,B)$ with $A=(a_{1},\dots,a_{p})$ and $B=(b_{1},\dots,b_{p})$, then $c(CS)=(a_{1},a_{2},\dots,a_{p})$.
@@ 3.4 Specialization to geometric sequences
Consider the geometric sequence $R=(a^{k},a^{k-1}b,\dots,b^{k})$ for $a.b,k\in\mathbb{N}$. Then $\gcd(R)=1$ if and only if $\gcd(a,b)=1$. When this occurs, we say $S=\langle R\rangle$ is a geometric numerical semigroup. As noted earlier, any geometric sequence is a special case of a compound sequence, which in turn is a special case of a free numerical semigroup. In the sequence $(a^{k},a^{k-1}b,\dots,b^{k})$, either all terms are odd, or all terms but one are even. By Corollary 3.7 and Corollary 3.6, we immediately have the following result.
@@ 3.5 Examples
Note that this can also be handled by Corollary 3.6 since all but one of the generating elements are even. Since the odd generating element is 55, the result is simply (551)/2(55-1)/2.
@@
Let $A=(2,3,4,3,3,3)$ and $B=(5,5,5,5,8,5)$. Then we have the compound sequence $C(A,B)$ and the numerical semigroup
@@ 4.1 Alternate approaches with even generating elements
First, numerical semigroups generated by compound sequences have a few easily describable Apéry sets. For instance, for $(A,B)$ a suitable pair and $C(A,B)=(n_{0},\dots,n_{p})$, there is an explicit description for the Apéry set of $S=\langle C(A,B)\rangle$ relative to any generating element $n_{i}$, as is given in [4, Theorem 15]:
@@
We should also mention that we used this description of $\operatorname{Ap}(S;n_{i})$ behind the scenes in Example 2.4. In that case, the numerical semigroup is $S=\langle 6,15,20\rangle$, generated by the compound sequence $C(A,B)$ where $A=(2,3)$ and B(5,4)B-(5,4). By the explicit description of $\operatorname{Ap}(S;n_{i})$ in Eq. (2),
@@
Alternatively, if $C(A,B)=(n_{0},\dots,n_{p})$ is a compound sequence and $n_{i}$ is some odd generating element, then we can reverse the order in which the terms $n_{0},\dots,n_{i}$ appear in the sequence (leaving the other terms where they are), to obtain a new sequence which generates the same numerical semigroup. This new sequence $T=(n_{i},n_{i-1},\dots,n_{0},n_{i+1},n_{i+2},\dots,n_{p})$ is not necessarily compound, but it is telescopic! (See [3, Proposition 3.10].) Then, for $A=(a_{1},\dots,a_{p})$ and $B=(b_{1},\dots,b_{p})$, we have $c(T)=(b_{i},\dots,b_{1},a_{i+1},\dots,a_{p})$. We now have a telescopic sequence with odd initial term, so we only need the result from Theorem 3.5 involving an odd initial term to compute O(G(S))E(G(S))O(G(S))-E(G(S)).

### v2

@@ Abstract
In this paper, we extend recent results about the distribution of even and odd gaps of a numerical semigroup. We find that, for any numerical semigroup, the distribution can be computed in terms of the numbers of or the sums of odd and even elements in a corresponding Apéry set. With free numerical semigroups specifically, we show that there are always at least as many odd gaps as even gaps, with equality precisely when the generating elements are all odd. We then specialize these results to the cases of numerical semigroups generated by compound and geometric sequences.
@@ 1 Introduction
In [undef], Cho, Lee, & Nam investigated the distribution of odd and even gaps of numerical semigroups. They gave general results in the case where one knows the Apéry set of a numerical semigroup relative to an odd generating element. For this, one needs only to know the difference between the numbers of even and odd elements of the Apéry set. They also gave explicit results for a few families of numerical semigroups which have Apéry sets that are known in the literature. Notably, in the case of a numerical semigroup $S$ generated by a compound sequence for which all minimal generating elements are odd, they showed that $S$ has as many odd gaps as it has even gaps. In their concluding remarks, they mentioned an open problem about computing the distribution of odd and even gaps in the case of a numerical semigroup generated by a compound sequence where at least one of the minimal generating elements is even.
@@ 1.1 Organization
In Section 3, we look specifically at the case of free numerical semigroups along with the sequences, known as telescopic sequences, that one uses to generate them. With knowledge of the Apéry sets of free numerical semigroups, we use an identity from Gassert & Shor [undefb] to compute the difference between the numbers of odd gaps and even gaps of any free numerical semigroup. Our main result is Theorem 3.5. Immediately afterward, we have two corollaries: with one we give a simple formula for the difference in the case where all but one of the generating elements of a free numerical semigroup are even; and with the other we show that a free numerical semigroup $S$ has the same number of odd gaps as even gaps precisely when the minimal generating elements of $S$ are all odd. We then apply these results to numerical semigroups generated by compound sequences and geometric sequences, and we round off this section with some explicit examples.
@@ 2.1 Preliminaries
For a comprehensive introduction to numerical semigroups, we refer the reader to [undefd].
@@
It is known that $\langle n_{1},\dots,n_{k}\rangle$ is a numerical semigroup if and only if $\gcd(n_{1},\dots,n_{k})=1$. Furthermore, every numerical semigroup $S$ can be written in the form $S=\langle n_{1},\dots,n_{k}\rangle$ for some sequence of positive integers $(n_{1},\dots,n_{k})$ with $\gcd(n_{1},\dots,n_{k})=1$. We say that $S$ is minimally generated by $(n_{1},\dots,n_{k})$ if no proper subsequence generates $S$. It happens that every numerical semigroup $S$ has a unique minimal generating sequence. Such a minimal generating sequence is necessarily finite.
@@
Put another way, $\operatorname{Ap}(S;t)$ consists of exactly $t$ elements, each the smallest element of $S$ in its congruence class modulo $t$. Many properties of a numerical semigroup can be derived directly from one of its Apéry sets. We will use the following identity from Gassert & Shor [undefb].
Proposition 2.2 ([undefb, Theorem 2.3]).
@@
where $\sigma(T)$ denotes the sum of the elements of a finite set $T$. (This result is originally due to Selmer. See [undefe, Section 2].)
@@ 2.2 Distribution of odd and even gaps given any Apéry set
Cho, Lee, & Nam give a formula ([undef, Theorem 3.1]) for O(G(S))E(G(S))O(G(S))-E(G(S)) in terms of $\operatorname{Ap}(S;t)$ in the case where $t$ is an odd minimal generating element of $S$. With Proposition 2.2, we are able to extend the result for any nonzero $t\in S$, odd or even. We also note that $t$ need not be a generating element of $S$ here.
@@
If $t$ is even, the function $f(n)=(-1)^{n}$ doesn’t help. We instead use $f(n)=n(-1)^{n}$ with Proposition 2.2:
@@
Since $t$ is even, $(-1)^{g+t}=(-1)^{g}$ and $\sum_{k=0}^{t-1}k(-1)^{k}=-t/2$. Thus,
@@
consisting of 16 odd gaps and 9 even gaps. We can verify directly that O(G(S))E(G(S))=169=7O(G(S))-E(G(S))=16-9=7.
@@ 3.1 Telescopic sequences and free numerical semigroups
We are now ready to specialize our results to free numerical semigroups. For a more detailed treatment, see [undeff] and the references therein.
@@
There are many families of sequences that are telescopic. For instance, any two-term $(a,b)$ sequence is telescopic, as is any geometric sequence $(a^{k},a^{k-1}b,\dots,b^{k})$. As we will see in Section 3.3, compound sequences are telescopic as well.
@@ 3.2 Distribution of odd and even gaps of a free numerical semigroup
$F_{j}=\left(\prod\limits_{\begin{subarray}{c}i=1,\\ i\neq j\end{subarray}}^{k}\sum\limits_{n_{i}=0}^{c_{i}-1}(-1)^{n_{i}t_{i}}\right)\cdot\sum\limits_{n_{j}=0}^{c_{j}-1}n_{j}t_{j}(-1)^{n_{j}t_{j}}.$
@@
Now, suppose $m>0$, which implies $\{0,\dots,m-1\}\subseteq I$. Since we have a positive number of even elements in $T$, we need to show the number of odd gaps of $S$ is strictly larger than the number of even gaps of $S$. Appealing to Theorem 3.5 once again, we have
In order to show $O(G(S))>E(G(S))$, it will suffice to show that
@@ 3.3 Specialization to compound sequences
We now focus on numerical semigroups generated by compound sequences, which were first described in [undefc].
@@
Note that the definition of a compound sequence in [undefc] specifies that $2\leq a_{i}<b_{i}$ for all $i$. With this condition in place, one has that $(n_{0},\dots,n_{p})$ is an increasing sequence. This is unnecessary for our work so we make no such restriction. The condition that $\gcd(a_{i},b_{j})=1$ for all $i\geq j$ is sufficient for our purposes.
The following proposition, from [undefa], says that the condition where $a_{i},b_{i}\geq 2$ for all $i$ gives rise to a minimal generating sequence.
@@
[undefa, Propositions 2.6 and 2.7] Suppose $(A,B)$ is a suitable pair with $C(A,B)=(n_{0},\dots,n_{p})$. Then $\gcd(n_{0},\dots,n_{p})=1$, implying that $\langle n_{0},\dots,n_{p}\rangle$ is a numerical semigroup. If we further have that $a_{i},b_{i}\geq 2$ for all $i$, then $(n_{0},\dots,n_{p})$ is a minimal generating sequence for this numerical semigroup.
@@
Proposition 3.10 ([undefb, Corollary 3.11]).
If $CS=(n_{0},\dots,n_{p})$ is a compound sequence, then $CS$ is a telescopic sequence. In particular, if $CS=C(A,B)$ for a suitable pair $(A,B)$ with $A=(a_{1},\dots,a_{p})$ and $B=(b_{1},\dots,b_{p})$, then $c(CS)=A=(a_{1},\dots,a_{p})$.
@@ 3.4 Specialization to geometric sequences
Consider the geometric sequence $R=(a^{k},a^{k-1}b,\dots,b^{k})$ for $a,b,k\in\mathbb{N}$. Then $\gcd(R)=1$ if and only if $\gcd(a,b)=1$. When this occurs, we say $S=\langle R\rangle$ is a geometric numerical semigroup. As noted earlier, any geometric sequence is a special case of a compound sequence, which in turn is a special case of a free numerical semigroup. In the sequence $(a^{k},a^{k-1}b,\dots,b^{k})$, either all terms are odd, or all terms but one are even. By Corollaries 3.6 and 3.7, we immediately have the following result.
@@ 3.5 Examples
Note that this can also be handled by Corollary 3.6 since $T$ is telescopic and all but one of the generating elements are even. Since the odd generating element is 55, the result is simply (551)/2(55-1)/2.
@@
Let $A=(2,3,4,3,3,3)$ and $B=(5,5,5,5,8,5)$, a suitable pair. We obtain the compound sequence $C(A,B)$ and corresponding numerical semigroup
@@ 4.1 Alternate approaches with even generating elements
First, numerical semigroups generated by compound sequences have a few easily describable Apéry sets. For instance, for $(A,B)$ a suitable pair and $C(A,B)=(n_{0},\dots,n_{p})$, there is an explicit description for the Apéry set of $S=\langle C(A,B)\rangle$ relative to any generating element $n_{i}$, as is given in [undefc, Theorem 15]:
@@
We should also mention that we used this description of $\operatorname{Ap}(S;n_{i})$ behind the scenes in Example 2.4. In that case, the numerical semigroup is $S=\langle 6,15,20\rangle$, generated by the compound sequence $C(A,B)$ where $A=(2,3)$ and $B=(5,4)$. By the explicit description of $\operatorname{Ap}(S;n_{i})$ in Eq. (2),
@@
Alternatively, if $C(A,B)=(n_{0},\dots,n_{p})$ is a compound sequence and $n_{i}$ is some odd generating element, then we can reverse the order in which the terms $n_{0},\dots,n_{i}$ appear in the sequence (leaving the other terms where they are), to obtain a new sequence which generates the same numerical semigroup. This new sequence $T=(n_{i},n_{i-1},\dots,n_{0},n_{i+1},n_{i+2},\dots,n_{p})$ is not necessarily compound, but it is telescopic! (See [undefb, Proposition 3.10].) Then, for $A=(a_{1},\dots,a_{p})$ and $B=(b_{1},\dots,b_{p})$, we have $c(T)=(b_{i},\dots,b_{1},a_{i+1},\dots,a_{p})$. We now have a telescopic sequence with odd initial term, so we only need the result from Theorem 3.5 involving an odd initial term to compute O(G(S))E(G(S))O(G(S))-E(G(S)).
@@ 4.2 Taking it further
Acknowledgments
We wish to thank Pieter Moree for the suggestion to look at [undef], which is how this paper came to be.